# Learned Free-Energy Functionals from Pair-Correlation Matching for Dynamical Density Functional Theory

### v2

@@ Abstract
Classical density functional theory (cDFT) and dynamical density functional theory (DDFT) are modern statistical mechanical theories for modeling many-body colloidal systems at the one-body density level. The theories hinge on knowing the excess free-energy accurately, which is however not feasible for most practical applications. Dijkman et al. [Phys. Rev. Lett. 134, 056103 (2025)] recently showed how a neural excess free-energy functional for cDFT can be learned from bulk simulations via pair-correlation matching. In this article, we demonstrate how this same functional can be applied to DDFT, without any retraining, to simulate non-equilibrium overdamped dynamics of inhomogeneous densities. We evaluate this on a 3D Lennard-Jones system with planar geometry under various complex external potentials and observe good agreement of the dynamical densities with those from expensive Brownian dynamic simulations, up to the limit of the adiabatic approximation. We further develop and apply an extension of DDFT based on gradient flows, to a grand-canonical system modeled after breakthrough gas adsorption studies, finding similarly good agreement. Our results demonstrate a practical route for leveraging learned free-energy functionals in DDFT, paving the way for accurate and efficient modeling of many-body non-equilibrium systems.
@@ I Introduction
Studying the collective behavior of many-body systems emerging from inter-particle interactions is fundamental in many diverse fields of science, ranging from gas adsorption studies in porous media [1] to semiconductors and protein adsorption [2]. Classical density functional theory (cDFT) [3] is a promising theory for accurately and efficiently modelling such many-body systems at the level of the one-body density profile $\rho(\mathbf{r})$ in thermodynamic equilibrium. cDFT relies on the existence of an excess free-energy functional $\mathcal{F}_{\text{exc}}[\rho(\mathbf{r})]$ which describes the inter-particle interactions, accounting for the non-ideal contribution to the total Helmholtz intrinsic free-energy functional $\mathcal{F}[\rho(\mathbf{r})]$. This functional is generally unknown and various approximations have been constructed ranging from the mean-field approximation [4] for soft interactions to fundamental measure theory [5] for hard-sphere interactions. More recently, neural functionals learned from Monte Carlo simulations have been proposed [6, 7, 8, 9, 10] and exhibit remarkable accuracy.
@@
In this article, we adopt the neural functional proposed by Dijkman et al. [10], that was trained by matching the pair-correlation function from bulk simulations, and evaluate it for DDFT. Despite only having seen bulk densities during training, we show that this functional produces accurate time evolution of $1$D inhomogeneous densities on a $3$D Lennard-Jones system with planar geometry. We further extend DDFT with a particle source term, based on gradient flows [17, 18], to model grand-canonical systems with non-conserved dynamics. We evaluate this extension on a system inspired by gas adsorption simulations, and observe good agreement with a hybrid Monte Carlo Brownian dynamic simulation.
@@ II Pair-correlation matching
Pair-correlation matching [10] is a method for learning an accurate neural excess free-energy functional $\mathcal{F}_{\theta,\text{exc}}$ (where $\theta$ represents neural network parameters) by training exclusively on a dataset of bulk radial distribution functions, rather than from heterogeneous density profiles. The method uses the relationship between the second functional derivative of the excess free-energy functional $\left(\delta^{2}\mathcal{F}_{\text{exc}}/\delta\rho(z_{i})\delta\rho(z_{j})\right)$ and the direct correlation function $c^{(2)}(\mathbf{r})$, which is related to the radial distribution function $g(r)$ via the Ornstein-Zernike equation ([4], Ch.3), for training. Specifically, a neural network $\mathcal{F}_{\theta,\text{exc}}$ is trained by optimizing a physics-based regularizer [19], that matches its Hessian $\left(\partial^{2}\mathcal{F}_{\text{exc}}/\partial\rho_{i}\partial\rho_{j}\right)$ with the direct correlation function obtained from Monte Carlo simulations of homogeneous bulk systems (Fig. 1). This functional is then used in the cDFT scheme to obtain accurate estimates for inhomogeneous equilibrium density profiles under various external fields without having used any inhomogeneous densities during training.
@@ III Classical dynamical density functional theory
DDFT was first introduced phenomenologically by Evans [3], and later microscopically derived from the Langevin equations by Marconi and Tarazona  [11], and from Smoluchowski equations by Archer and Evans [12]. An extensive review of DDFT fundamentals can be found in [13]. DDFT has the form of a continuity equation where $\partial_{t}\rho(\mathbf{r},t)$ is proportional to the divergence of a vector field as follows,
@@
where $\mathcal{F}_{\text{exc}}([\rho])$ is the excess free-energy functional, $V_{\text{ext}}$ is the external potential, $\beta=\left(1/k_{B}T\right)$, and $\Lambda$ is the thermal wavelength. Expanding eq. 1 further we get,
@@
In the ideal gas limit, where there are no inter-particle interactions i.e. $\mathcal{F}_{\text{exc}}[\rho]=0$, this equation reduces to a drift-diffusion equation where particles only diffuse and respond to an external potential. The DDFT equation in this standard form, with no velocity or inertial terms, describes purely overdamped dynamics where frictional forces dominate. It also relies on the adiabatic approximation which assumes that the non-equilibrium pair correlation function is equivalent to its equilibrium counterpart at the same one-body density. However, for systems where superadiabatic forces such as memory effects are significant, more advanced frameworks such as power functional theory [20, 21] and superadiabatic DDFT [22] are required. It is also worth noting that since standard DDFT has the form of a continuity equation and conserves the particle number (neglecting the boundaries), it is only a canonical theory in this regard. In practice however it is common to still adopt the same grand-canonical free energy functional from cDFT under the adiabatic approximation. We follow this practice and adopt the equilibrium neural excess-free energy functional from Dijkman et al. [10].
@@ IV Extension to open systems
Many practical systems are open to particle exchange with a reservoir—for instance, gas adsorption in porous materials is often modeled using hybrid grand-canonical molecular dynamics simulations [23, 24]. We develop a novel extension of DDFT for such hybrid setups, allowing for particle insertion and removal through a reservoir coupling. This setting was studied before by Thiele et al. [25], where a source term was added to model the evaporation of a colloidal suspension into a vapor reservoir. We develop a similar extension, but from a principled Wasserstein-Fisher-Rao gradient flow perspective.
The 2-Wasserstein distance between two densities $\rho_{0},\rho_{1}$, defined as
[…]
where $\Gamma(\rho_{0},\rho_{1})$ is the space of all joint distributions with marginals $\rho_{0}$ and $\rho_{1}$, can also be expressed in dynamical form
[…]
where $V(\rho_{0},\rho_{1})$ is the set of velocity fields $v_{t}$ such that $\left(\partial\rho_{t}/\partial t\right)=-\nabla_{\mathbf{r}}.(\rho_{t}(\mathbf{r})v_{t}(\mathbf{r}))$ with $\rho_{t=0}=\rho_{0}$ and $\rho_{t=1}=\rho_{1}$. From this dynamical formulation, which was first proposed by Benamou & Brenier [26], it can be shown that the flow defined by DDFT (eq. 1) is in fact the steepest descent on the free-energy functional $\mathcal{F}[\rho]$ in the space of densities under the 2-Wasserstein distance metric [27].
The space of densities can be endowed with different metrics [17] which correspond to different optimization dynamics. In particular, the Wasserstein Fisher-Rao distance [28, 18] can be defined by extending  eq. 5 with a growth term,
[…]
and the corresponding continuity equation is updated as $\left(\partial\rho_{t}/\partial t\right)=-\nabla_{\mathbf{r}}.(\rho_{t}(\mathbf{r})v_{t}(\mathbf{r}))+g_{t}(\mathbf{r})\rho_{t}(\mathbf{r})$. It can be shown that the flow that minimizes the free-energy functional $\mathcal{F}[\rho]$ under this Wasserstein Fisher-Rao metric is given by the following PDE (proof in [18]):
@@
Figure 2: Time evolution of the one-body density profile $\rho(z,\tau)$ for a system of interacting $3$D LJ particles with planar geometry subjected to a repulsive Gaussian potential $V_{\text{ext}}(z)$ (grey curve, right axis). The plots compare two neural excess free-energy functionals and an analytical FMT functional, with a reference Brownian dynamics simulation. Initially ($t=0\tau$), the system is at equilibrium with a flat density profile. Upon application of the external potential, the density profile evolves non-trivially, exhibiting a depletion near the potential maximum and eventually approaching a new equilibrium. The neural DDFT approaches (red and blue) exhibit good agreement with Brownian dynamics throughout the temporal evolution, notably with the pair-correlation matching approach (red) not having seen any inhomogeneous densities during training.
While standard DDFT (eq. 1) defines the change of density along a vector field $v_{t}(\mathbf{r})=\nabla_{\mathbf{r}}\left(\delta\mathcal{F}[\rho]/{\delta\rho_{t}}\right)$, this PDE also defines the addition and removal of density with a growth term that is proportional to the difference between the local chemical potential $\left({\delta\mathcal{F}\left[\rho_{t}\right]}/{\delta\rho_{t}}(\mathbf{r})\right)$ and the average chemical potential of the whole system. We replace this average potential with a target chemical potential $\mu$ to model a grand canonical reservoir and get the following form,
$\begin{split}\frac{\partial\rho(\mathbf{r},t)}{\partial t}=\,&\underbrace{\Gamma\nabla\cdot\Bigg{[}\rho(\mathbf{r},t)\nabla\Bigg{(}\frac{\delta\mathcal{F}[\rho]}{\delta\rho(\mathbf{r},t)}\Bigg{)}\Bigg{]}}_{\text{continuity equation}}\\ &+\underbrace{(1-\lambda(\mathbf{r}))\Gamma^{\prime}\Bigg{(}\mu-\frac{\delta\mathcal{F}[\rho]}{\delta\rho(\mathbf{r},t)}\Bigg{)}\rho(\mathbf{r},t)}_{\text{growth term}},\end{split}$ (8)
@@
$\mu=V_{\text{ext}}+\frac{1}{\beta}\ln{\left(\Lambda^{3}\rho_{\text{eq}}(\mathbf{r})\right)}+\left.\frac{\delta\mathcal{F}_{\text{exc}}[\rho(\mathbf{r})]}{\delta\rho(\mathbf{r})}\right|_{\rho=\rho_{\text{eq}}}.$ (9)
This relation suggests that the contribution from the external potential $V_{\text{ext}}$, ideal gas $\ln\rho_{\text{eq}}(\Lambda^{3}\mathbf{r})$, and the functional derivative of the excess free energy add up at each position to $\mu$ at equilibrium. Using this conservation relation we can write the growth-term PDE as,
$\begin{split}\frac{\partial\rho(\mathbf{r},t)}{\partial t}=-\Bigg{(}&V_{\text{ext}}+\frac{1}{\beta}\ln{\left(\Lambda^{3}\rho(\mathbf{r},t)\right)}\\ &+\frac{\delta\mathcal{F}_{\text{exc}}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}-\mu\Bigg{)}\rho(\mathbf{r},t).\end{split}$ (10)
@@
$\left\|\nabla\frac{\delta\mathcal{F}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}\right\|=0\iff\frac{\delta\mathcal{F}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}=\mu,$ (11)
@@ V Implementation
Figure 3: The final equilibrium density profiles and their RMSE over time evolution under complex external potentials (see Appendix E for the movie version). (a)-(d) A comparison of the final 1D equilibrium density profiles for a three-dimensional LJ system with planar geometry under different external potentials constructed from sums of repulsive Gaussian and steep well potentials. (e)-(h) show the corresponding RMSE deviation from the Brownian dynamics results over time evolution. The densities obtained with the neural functionals are significantly more accurate than those with the analytical functional, notably with the pair-correlation matching approach (red) not having seen any inhomogeneous densities during training.

### v3

@@ Abstract
Classical density functional theory (cDFT) and dynamical density functional theory (DDFT) are modern statistical mechanical theories for modeling many-body colloidal systems at the one-body density level. The theories hinge on knowing the excess free-energy accurately, which is however not feasible for most practical applications. Dijkman et al. [Phys. Rev. Lett. 134, 056103 (2025)] recently showed how a neural excess free-energy functional for cDFT can be learned from bulk simulations via pair-correlation matching. In this article, we demonstrate how this same functional can be applied to DDFT, without any retraining, to simulate non-equilibrium overdamped dynamics of inhomogeneous densities. We evaluate this on a three-dimensional Lennard-Jones system with planar geometry under various complex external potentials and observe good agreement of the dynamical densities with those from expensive Brownian dynamic simulations, up to the limit of the adiabatic approximation. We further develop and apply an extension of DDFT based on gradient flows, to a grand-canonical system modeled after breakthrough gas adsorption studies, finding similarly good agreement. Our results demonstrate a practical route for leveraging learned free-energy functionals in DDFT, paving the way for accurate and efficient modeling of many-body non-equilibrium systems.
@@ I Introduction
Studying the collective behavior of many-body systems emerging from inter-particle interactions is fundamental in many diverse fields of science, ranging from gas adsorption studies in porous media [1] to semiconductors and protein adsorption [2]. Classical density functional theory (cDFT) [3] is a promising theory for accurately and efficiently modeling such many-body systems at the level of the one-body density profile $\rho(\mathbf{r})$ in thermodynamic equilibrium. cDFT relies on the existence of an excess free-energy functional $\mathcal{F}_{\text{exc}}[\rho(\mathbf{r})]$ which describes the inter-particle interactions, accounting for the non-ideal contribution to the total Helmholtz intrinsic free-energy functional $\mathcal{F}[\rho(\mathbf{r})]$. This functional is generally unknown and various approximations have been constructed ranging from the mean-field approximation [4] for soft interactions to fundamental measure theory [5] for hard-sphere interactions. More recently, neural functionals learned from Monte Carlo simulations have been proposed [6, 7, 8, 9, 10] and exhibit remarkable accuracy.
@@
In this article, we adopt the neural functional proposed by Dijkman et al. [10], which was trained by matching the pair-correlation function from bulk simulations, and we evaluate it for DDFT. Despite only having seen bulk densities during training, we show that this functional produces accurate time evolution of 11D inhomogeneous densities on a 33D Lennard-Jones system with planar geometry. We further extend DDFT with a particle source term, based on gradient flows [17, 18], to model grand-canonical systems with non-conserved dynamics. We evaluate this extension on a system inspired by gas adsorption simulations, and observe good agreement with a hybrid Monte Carlo Brownian dynamic simulation.
@@ II Pair-correlation matching
Pair-correlation matching [10] is a method for learning an accurate neural excess free-energy functional $\mathcal{F}_{\theta,\text{exc}}$ (where $\theta$ represents neural network parameters) by training exclusively on a dataset of bulk radial distribution functions, rather than from heterogeneous density profiles. The method uses the relationship between the second functional derivative of the excess free-energy functional $\left(\delta^{2}\mathcal{F}_{\text{exc}}/\delta\rho(z_{i})\delta\rho(z_{j})\right)$ and the direct correlation function $c^{(2)}(\mathbf{r})$, which is related to the radial distribution function $g(r)$ via the Ornstein-Zernike equation ([4], Chap.3), for training. Specifically, a neural network $\mathcal{F}_{\theta,\text{exc}}$ is trained by optimizing a physics-based regularizer [19], which matches its Hessian $\left(\partial^{2}\mathcal{F}_{\text{exc}}/\partial\rho_{i}\partial\rho_{j}\right)$ with the direct correlation function obtained from Monte Carlo simulations of homogeneous bulk systems (Fig. 1). This functional is then used in the cDFT scheme to obtain accurate estimates for inhomogeneous equilibrium density profiles under various external fields without having used any inhomogeneous densities during training. In addition to pair-correlation matching, we can also train a neural free-energy functional using single-body direct correlation matching; this is further detailed with a brief recapitulation of the key cDFT equations in Appendix A.
@@ III Classical dynamical density functional theory
DDFT was first introduced phenomenologically by Evans [3], and later microscopically derived from the Langevin equations by Marconi and Tarazona  [11], and from Smoluchowski equations by Archer and Evans [12]. An extensive review of DDFT fundamentals can be found in [13]; see Appendix B for a concise derivation. DDFT has the form of a continuity equation where $\partial_{t}\rho(\mathbf{r},t)$ is proportional to the divergence of a vector field as follows:
@@
where $\mathcal{F}_{\text{exc}}([\rho])$ is the excess free-energy functional, $V_{\text{ext}}(\mathbf{r})$ is the external potential, $\beta=\left(1/k_{B}T\right)$, and $\Lambda$ is the thermal wavelength. Expanding eq. 1 further we get,
@@
In the ideal gas limit, where there are no inter-particle interactions i.e. $\mathcal{F}_{\text{exc}}[\rho]=0$, this equation reduces to a drift-diffusion equation where particles only diffuse and respond to an external potential. The DDFT equation in this standard form, with no velocity or inertial terms, describes purely overdamped dynamics where frictional forces dominate. It also relies on the adiabatic approximation which assumes that the non-equilibrium pair correlation function is equivalent to its equilibrium counterpart at the same one-body density. However, for systems where superadiabatic components such as rotational flows or memory effects are significant, more advanced frameworks such as power functional theory [20, 21] and superadiabatic DDFT [22] are required. It is also worth noting that since standard DDFT has the form of a continuity equation and conserves the particle number (neglecting the boundaries), it is only a canonical theory in this regard. In practice however it is common to still adopt the same grand-canonical free energy functional from cDFT under the adiabatic approximation. We follow this practice and adopt the equilibrium neural excess-free energy functional from Dijkman et al[10].
@@ IV Extension to open systems
Many practical systems are open to particle exchange with a reservoir—for instance, gas adsorption processes in porous materials—and these are often modeled using hybrid grand-canonical molecular dynamics simulations [23, 24]. We develop a novel extension of DDFT for such hybrid setups, allowing for particle insertion and removal through a reservoir coupling. This setting was studied before by Thiele et al. [25], where a source term was added to model the evaporation of a colloidal suspension into a vapor reservoir. We develop a similar extension, but from a principled Wasserstein-Fisher-Rao gradient flow perspective.
The 2-Wasserstein distance between two densities $\rho_{0},\rho_{1}$ is defined as
$\operatorname{W}_{2}^{2}(\rho_{0},\rho_{1})=\inf_{\rho\in\Gamma(\rho_{0},\rho_{1})}\int\rho(\mathbf{r}_{1},\mathbf{r}_{2})\|\mathbf{r}_{1}-\mathbf{r}_{2}\|_{2}^{2}d\mathbf{r}_{1}d\mathbf{r}_{2},$ (4)
where $\Gamma(\rho_{0},\rho_{1})$ is the space of all joint distributions with marginals $\rho_{0}$ and $\rho_{1}$. The infima gives the optimal coupling between the densities $\rho_{0}$, $\rho_{1}$ that minimizes their total expected Euclidean distance. This can also be expressed in dynamical form as
$\operatorname{W}_{2}^{2}(\rho_{0},\rho_{1})=\inf_{v_{t}\in V(\rho_{0},\rho_{1}),\rho_{t}}\int_{0}^{1}\mathbb{E}_{\rho_{t}(\mathbf{r})}[\|v_{t}(\mathbf{r})\|_{2}^{2}]dt,$ (5)
where $V(\rho_{0},\rho_{1})$ is the set of velocity fields $v_{t}$ such that $\left(\partial\rho_{t}/\partial t\right)=-\nabla_{\mathbf{r}}.(\rho_{t}(\mathbf{r})v_{t}(\mathbf{r}))$ with $\rho_{t=0}=\rho_{0}$ and $\rho_{t=1}=\rho_{1}$. In other words, the 2-Wasserstein distance is the optimal $(v_{t},\rho_{t})$ pair, satisfying the continuity equation, that minimizes the total expected kinetic energy. From this dynamical formulation, which was first proposed by Benamou and Brenier [26], it can be shown that the flow defined by DDFT [eq. 1] is in fact the steepest descent on the free-energy functional $\mathcal{F}[\rho]$ in the space of densities under the 2-Wasserstein distance metric [27].
The space of densities can be endowed with different metrics [17] which correspond to different optimization dynamics. In particular, the Wasserstein Fisher-Rao distance [28, 18] can be defined by extending  eq. 5,
$\operatorname{WFR}^{2}(\rho_{0},\rho_{1})=\inf_{v_{t},\rho_{t},g_{t}}\int_{0}^{1}\mathbb{E}_{\rho_{t}(\mathbf{r})}[\|v_{t}(\mathbf{r})\|_{2}^{2}+\lambda g_{t}(\mathbf{r})^{2}]dt,$ (6)
where $g_{t}$ is a growth term controlling the addition and removal of density, and $\lambda$ is a constant balancing the cost of the flow and growth terms. The corresponding continuity equation is updated as $\left(\partial\rho_{t}/\partial t\right)=-\nabla_{\mathbf{r}}.(\rho_{t}(\mathbf{r})v_{t}(\mathbf{r}))+g_{t}(\mathbf{r})\rho_{t}(\mathbf{r})$. Thus, the WFR distance allows for unbalanced transport where the total particle number can change. It can be shown that the flow that minimizes the free-energy functional $\mathcal{F}[\rho]$ under this Wasserstein Fisher-Rao metric is given by the following PDE (proof in [18]):
@@
Figure 2: Time evolution of the one-body density profile $\rho(z,\tau)$ for a system of interacting 33D LJ particles with planar geometry subjected to a repulsive Gaussian potential $V_{\text{ext}}(z)$ (grey curve, right axis). The plots compare two neural excess free-energy functionals and an analytical FMT functional, with a reference Brownian dynamics simulation. Initially ($t=0\tau$), the system is at equilibrium with a flat density profile. Upon application of the external potential, the density profile evolves non-trivially, exhibiting a depletion near the potential maximum and eventually approaching a new equilibrium ($t=3\tau$). The neural DDFT approaches (red and blue) exhibit good agreement with Brownian dynamics throughout the temporal evolution, notably with the pair-correlation matching approach (red) not having seen any inhomogeneous densities during training.
While standard DDFT [eq. 1] defines the change of density along a vector field $v_{t}(\mathbf{r})=\nabla_{\mathbf{r}}\left(\delta\mathcal{F}[\rho]/{\delta\rho_{t}}\right)$, this PDE also defines the addition and removal of density with a growth term that is proportional to the difference between the local chemical potential $\left({\delta\mathcal{F}\left[\rho_{t}\right]}/{\delta\rho_{t}}(\mathbf{r})\right)$ and the average chemical potential of the whole system. We replace this average potential with a target chemical potential $\mu_{\text{T}}$ to model a grand canonical reservoir and get the following form,
$\begin{split}\frac{\partial\rho(\mathbf{r},t)}{\partial t}=\,&\underbrace{\Gamma\nabla\cdot\Bigg[\rho(\mathbf{r},t)\nabla\Bigg(\frac{\delta\mathcal{F}[\rho]}{\delta\rho(\mathbf{r},t)}\Bigg)\Bigg]}_{\text{continuity equation}}\\ &+\underbrace{(1-\lambda(\mathbf{r}))\Gamma^{\prime}\Bigg(\mu_{\text{T}}-\frac{\delta\mathcal{F}[\rho]}{\delta\rho(\mathbf{r},t)}\Bigg)\rho(\mathbf{r},t)}_{\text{growth term}},\end{split}$ (8)
@@
$\mu_{\text{eq}}=V_{\text{ext}}(\mathbf{r})+\frac{1}{\beta}\ln{\left(\Lambda^{3}\rho_{\text{eq}}(\mathbf{r})\right)}+\left.\frac{\delta\mathcal{F}_{\text{exc}}[\rho(\mathbf{r})]}{\delta\rho(\mathbf{r})}\right|_{\rho=\rho_{\text{eq}}}.$ (9)
This relation shows that at equilibrium the contribution from the external potential $V_{\text{ext}}(\mathbf{r})$, ideal gas $\ln\left(\Lambda^{3}\rho_{\text{eq}}(\mathbf{r})\right)$, and the functional derivative of the excess free energy add up at each position to $\mu_{\text{eq}}$. Using this conservation relation we can write the growth-term PDE as,
$\begin{split}\left.\frac{\partial\rho(\mathbf{r},t)}{\partial t}\right|_{\text{gr.}}=-\Gamma^{\prime}\Bigg(&V_{\text{ext}}(\mathbf{r})+\frac{1}{\beta}\ln{\left(\Lambda^{3}\rho(\mathbf{r},t)\right)}\\ &+\frac{\delta\mathcal{F}_{\text{exc}}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}-\mu_{\text{T}}\Bigg)\rho(\mathbf{r},t).\end{split}$ (10)
@@
$\left\|\nabla\frac{\delta\mathcal{F}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}\right\|=0\iff\frac{\delta\mathcal{F}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}=\mu_{\text{T}},$ (11)
@@ V Implementation
Figure 3: The final equilibrium density profiles and their RMSE over time evolution under complex external potentials (see SM [29] for the movie version). (a)-(d) A comparison of the final 1D equilibrium density profiles for a three-dimensional LJ system with planar geometry under different external potentials constructed from sums of repulsive Gaussian and steep well potentials. (e)-(h) show the corresponding RMSE deviation from the Brownian dynamics results over time evolution. The densities obtained with the neural functionals are significantly more accurate than those with the analytical functional, notably with the pair-correlation matching approach (red) not having seen any inhomogeneous densities during training.